OpenStudy (anonymous):

Calculate following derivatives: \[\Large \frac{d}{dx} \int_0^{x^{2}} dt \frac{\exp t}{t^{2}+1}\]

OpenStudy (anonymous):

\[ 2x\frac {e^{x^2}}{x^4+1} \]

OpenStudy (anonymous):

FTC and Chain Rule

OpenStudy (anonymous):

Mr Saab is it end solution so that i can write it tomorrow on exam if same question appears?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \frac {d}{dx}\int_{v(x)}^{u(x)} f(x)dx = f(u(x)) u'(x)- f(v(x)) v'(x) \]

OpenStudy (anonymous):

In your case \[ v(x)=0\\ u(x)=x^2 \]

OpenStudy (anonymous):

ok thank you very very much mr Saab

OpenStudy (anonymous):

\[\frac {d}{dx}\int_{v(x)}^{u(x)} f(t)dt = f(u(x)) u'(x)- f(v(x)) v'(x)\] A small misprint

OpenStudy (anonymous):

ok thank you